OpenStudy (anonymous):

adanne1990's question

OpenStudy (anonymous):

Can someone solve these problems (calculus 2 I believe). If not then I'll try to do them myself, but it's just that im tryin to do my hw..

OpenStudy (anonymous):

u want them all solved?

OpenStudy (anonymous):

not me someone else does.. but im workin on my hw.. and if anyone feels like solving a couple of problems then go ahead.. adanne1990 emailed them to me askin for help.

OpenStudy (anonymous):

o, ok

OpenStudy (amistre64):

If it didnt say "TEST" right there on it, id be more willing to oblige ... :/

OpenStudy (anonymous):

lol want me to rename the file?

OpenStudy (amistre64):

and rewrite it as well ;)

OpenStudy (anonymous):

sure..

OpenStudy (anonymous):

lolz

OpenStudy (anonymous):

here are a couple of problems..

OpenStudy (amistre64):

this is the drawing ... not to scale of course

OpenStudy (amistre64):

your integration for area will be from 1 to 5; and from 5 to 9 \[\int_{1}^{5}(x+1)-(2)\ dx+\int_{5}^{9}(11-x)-(2)\ dx\] if I see it right

OpenStudy (amistre64):

convert the equations to inverses: y = x+1 ; x = y-1 y =11-x ; x = 11-y and integrate with respect to y from 2 to 7 \[\int_{2}^{7}(y-1)-(11-y)\ dy\]

OpenStudy (amistre64):

lol ... typoes it :)

OpenStudy (amistre64):

\[\int_{}^{}(y-1)-(11-y)\ dy\] .. nah its good; gonna put tha tpersonaility in a time out :)

OpenStudy (anonymous):

thanks man.. mind finishin it off?

OpenStudy (amistre64):

y-1-11+y = 2y-12 {S} 2y-12 dy --> (2y^2)/2 -12y; y^2 -12y area = (7)^2-12(7)-[(2)^2-12(2)] = 49 - 84 -(4-24) = -35 -(-20) = -15 just caught a fauxpaux; the point above needs to be (5,6); 5+1 = 6; 11 - 5 = 6 .. which alters the integrals a little and means this set up is errored. lets try that with y from 2 to 6 area = (6)^2-12(6)-[(2)^2-12(2)] = 36 - 72 -(4-24) = -36 -(-20) = -16 <-- this negative here tells me that I just subtracted the function in the wrong order; well, not wrong, just backwards frm normality :) ill revamp this for accuracy

OpenStudy (amistre64):

A). graph is good except need the points are: \begin{array}c &(5,6)&\\ (1,2)&...&(9,2)\\ \end{array} B). \(\int_{1}^{5}(x+1)-(2)dx+\int_{5}^{9}(11-x)-(2)dx\) C). \(\int_{2}^{6}(11-y)-(y-1)dy\) D). \(F(y)=\int_{2}^{6}(11-y)-(y-1)dy\) \(=\int_{2}^{6}(11-y-y+1)dy\) \(=\int_{2}^{6}(12-2y)dy\) \(=12y-y^2\) \begin{align} Area &= 12(6)-(6)^2-(12(2)-(2)^2)\\ &=72-36-(24-4)\\ &=36-20\\ &=16 \end{align}

OpenStudy (amistre64):

when we rotate about the y axis; we get a radius equal to x and a height equal to f(x); so our integration goes from 1 to 9; but becasue of the change in f(x) at 5 we need to split it up; \[2\pi\left(\int_{1}^{5}x[(x+1)-2]dx+\int_{5}^{9}x[(11-x)-2]dx\right)\]

OpenStudy (amistre64):

x(x-1) = x^2 - x {S} x^2 - x dx --> x^3/3 -x^2/2 5^3/3 -5^2/2 - (1^3/3 -1^2/2) 125/3 -25/2 - 1/3 -1/2 124/3 -24/2 124(2)/6 -24(3)/6 248/6 -72/6 176/6 = 29' 1/3 x(9-x) = 9x - x^2 {S} 9x - x^2 dx --> 9x^2/2 -x^3/3 9(9)^2/2 -(9)^3/3 - (9(5)^2/2 -(5)^3/3) 729/2 -729/3 - 225/2 +125/3 504/2 -604/3 504(3)/6 -604(2)/6 1512/6 -1208/6 304/6 = 50' 2/3 50' 2/3 + 29' 1/3 --------- 79' 3/3 = 80 2pi(80) = 160pi for the volume of revolution

OpenStudy (amistre64):

its best to dbl chk my work, I am prone to errors in this :)

OpenStudy (amistre64):

but thats about me limit for this; good luck with it :)

OpenStudy (anonymous):

hey assassin dou u have a facebook?